# PlantTraitNet: An Uncertainty-Aware Multimodal Framework for Global-Scale Plant Trait Inference from Citizen Science Data

### v1

@@ Abstract
Extended versionhttps://arxiv.org/abs/¡¿
@@ Introduction
Terrestrial plants, as the largest primary producers on Earth, contribute about 60% to the global net primary productivity (field1998primary) and play a critical role in the carbon and energy cycles of our Earth system (pan2011large; schlesinger2020biogeochemistry). However, understanding how plants influence these cycles is challenging, as the functioning of plants varies profoundly according to their traits. For instance, traits such as canopy height and leaf area control resource acquisition, while leaf tissue properties, such as nitrogen content or dry matter content, are indicators of plant resilience (diaz2016global). Although these traits are essential for understanding ecosystem processes, the data on such traits is sparse, as their measurement involves costly field surveys and laboratory analysis. Global plant trait databases such as TRY (kattge2011try) aggregate thousands of trait measurements from numerous studies and regions, providing an invaluable resource for functional biogeography and ecosystem modelling. However, even with these collective efforts, significant gaps persist in the geographic coverage of trait data across biomes, ecosystems, and species, constraining our ability to fully understand and predict global patterns of vegetation function and change  (diaz2016global; kattge2020try).
@@
Given the strong link between plant morphology and function, plant photographs in concert with computer vision offer a promising avenue for large-scale estimation of plant traits. Citizen science platforms such as iNaturalist (su2021semi) and Pl@ntNet (garcin2021plantnet) have collected more than 50 million research-grade plant photographs around the world, creating a unique resource for uncovering global plant trait distributions (wolf2022citizen). These datasets, primarily curated for the identification of plant species, provide plant images and species labels, but do not provide trait annotations  (plantclef-2025; stevens2024bioclip; van2018inaturalist). However, prior work has demonstrated that trait information for these species can be indirectly obtained by linking species names from citizen science records to trait databases such as TRY (schiller2021deep; wolf2022citizen). Through this species-level matching, trait values can be weakly assigned to images, enabling the construction of large-scale, trait-annotated image datasets. These datasets can then be used to train scalable computer vision models for trait prediction from images (schiller2021deep). Such models enable direct trait estimation from photographs, independent of whether the species is known or if a record exists in a trait database. Here, we attempt to advance this approach by predicting multiple traits simultaneously, leveraging shared visual features and underlying trait correlations. Subsequently, we spatially aggregate trait predictions derived from individual geotagged photographs to create global, gridded geospatial maps representing the trait distributions across plant communities and ecosystems  (schiller2021deep).
@@
The geolocation of each photograph not only enables the spatial aggregation of the predictions into geospatial maps but also allows for the integration of spatial context into the prediction process itself  (schiller2021deep). For instance, climate, including temperature and precipitation, or phenological information from satellite data is known to be the key variable shaping global trait distributions, making it a promising predictor  (bruelheide2018global; schiller2021deep; joswig2022climatic). However, integrating large-scale geospatial products can be challenging due to data gaps and feature selection (lusk2025smartphones).
@@
Recent advances in geospatial foundation models (GeoFMs) now support the seamless integration of such context into downstream tasks. Examples include Climplicit (dollinger2025climplicit), which encodes climate information, and SatCLIP (klemmer2025satclip), which leverages satellite Earth observation data. Such GeoFMs have demonstrated strong generalization in global mapping applications. In this study, we test the integration of such GeoFMs into visual trait prediction to enhance performance through geospatial context.
@@
An important challenge for computer vision with citizen science data is its inherent heterogeneity and noise (sierra2024divshift), ranging from inconsistent image quality due to varied photo acquisition methods (feature noise) to ambiguous trait annotations from weak supervision (label noise). Such data characteristics may result in implausible predictions and leave an imprint on the aggregation into global trait products. Moreover, such training data noise can substantially degrade a model’s ability to generalize to unseen data  (lu2022selc; arpit2017closer; zhang2021understanding). To overcome both feature and label noise in the citizen-science data, we propose an uncertainty-aware probabilistic deep learning framework that estimates predictive uncertainty. The predicted uncertainty is used to dynamically down-weight highly noisy samples and to filter out unreliable data points, thereby reducing overfitting to spurious patterns.
@@
  • We apply PlantTraitNet on more than 300K independent samples of citizen science photos and spatailly aggreagate the predictions to global trait maps. A benchmark against globally distributed vegetation survey data (sPlotOpen) revealed that these PlantTraitNet-derived traits maps consistently outperform previous global trait products.
@@
Pioneering work by schiller2021deep showed that plant traits, such as height, nitrogen content, specific leaf area, or leaf area, can be predicted from citizen science images using weak supervision, where species-level trait labels are derived from the TRY database (kattge2020try). While  schiller2021deep focused on single-task models, (cherif2023spectra) showed that predicting multiple plant traits simultaneously can exploit trait-trait correlations and joint features in the predictor data.
However, schiller2021deep did not assess whether weak supervision enables capturing within-species trait variation (e.g., size differences among individuals of the same species). Moreover,  schiller2021deep did not test how aggregating individual predictions on a global scale resembles large-scale trait variation across the biosphere. wolf2022citizen provided an approach to validate global trait maps using vegetation survey data of plant communities from the collaborative initiative sPlot (bruelheide2018global; sabatini2021splotopen) linked with trait data from the TRY database (kattge2020try). This approach provides an effective means to evaluate the potential of computer vision models for generating trait maps at global scale.
A persistent challenge with citizen science data is the noise in both images and labels (sierra2024divshift; schiller2021deep), often structured spatially. Such noise can bias both inference and training, as deep networks tend to memorize noisy labels, compromising generalization (lu2022selc; arpit2017closer; zhang2021understanding).
@@
  • Using visual and depth-based foundation models (oquab2023dinov2; yang2024depth) to better represent heterogeneous plant imagery.
  • Leveraging multi-task learning to exploit trait correlations (cherif2023spectra).
  • Incorporating uncertainty-aware training to address label noise (yeo2021robustness; jiang2024uncertainty).
  • Benchmarking global trait predictions against sPlot vegetation survey data (wolf2022citizen).
@@
Figure 1: Geographic coverage of the citizen science data (top) and independent benchmark reference data (bottom) from vegetation surveys (sPlotOpen, (sabatini2021splotopen)).
@@
To predict plant traits at a global scale, we utilize two large-scale citizen science datasets: iNaturalist  (gbif2025; su2021semi) and Pl@ntNet-300K (garcin2021plantnet). These datasets consist of plant images annotated with species labels and geolocations but lack direct trait measurements. Following schiller2021deep, we weakly annotate each image using species-level trait distributions from the TRY database (kattge2020try), based on the premise that interspecific trait variation generally exceeds intraspecific variation (dong2020components; wright2017global).
We model each trait as a normal distribution per species, using TRY-derived means and standard deviations, and sample trait values within the interquartile range to reduce outlier influence. To account for intraspecific variability, we resample traits for each image at every training epoch  (schiller2021deep).
@@
We further reduce noise through model-driven uncertainty estimates (see Methodology). The final data includes  230K training images across  5K species and  80K images in validation set.
@@
For evaluation, we use the sPlotOpen database (sabatini2021splotopen). The georeferenced sPlotOpen records represent plant community compositions, which were linked with trait data from the TRY database (kattge2020try). This data provides global trait maps of community-weighted mean (CWM) trait values.
@@
Figure 4: Overview of the pipeline. We filter weakly labeled citizen science data (Raw data) based on high model uncertainty (Step 1) and large residuals from species trait medians (Step 2). We use this refined data for training the models (Step 3), which are evaluated by comparing spatially aggregated predictions (1° resolution) against overlapping vegetation surveys (sPlotOpen)
@@
Given an input image $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$, we use the pretrained DINOv2 ViT-B/14 encoder to extract a sequence of patch-level feature embeddings $\in\mathbb{R}^{N\times C}$  (oquab2023dinov2).
@@
A novel addition to our architecture is the use of depth priors from foundation models for monocular depth estimation. While standard 2D RGB images lack explicit three-dimensional spatial cues, depth information encodes the distance between the sensor and surface points on the plant, enabling a more accurate reconstruction of plant morphology and structure. To incorporate depth, we use the pretrained and frozen encoder from the Depth-Anything-v2(DA-V2) model  (yang2024depth), denoted as $h$. Although various models could be used, we adopt DA-V2 for its strong generalization capabilities, attributed to its training on large-scale labeled and unlabeled datasets and its student-teacher distillation framework. We use the ViT-B variant, which outputs a set of embeddings $h(I)\in\mathbb{R}^{N\times C}$. Similar to the image encoder, we apply adaptive average pooling along the patch token dimension $N$, reducing it to a 64-dimensional representation. This pooled output is flattened and passed through a MLP to generate the depth prior embedding of dimension 768.
@@
Plants are tailored to local climatic conditions, such as precipitation and temperature, through their traits (joswig2022climatic).
To incoporate this climatic context as cue in the prediction process, we integrate Climplicit (dollinger2025climplicit) into our architecture, a spatio-temporal geo-location encoder trained on the CHELSA climate dataset (karger2017climatologies). Climplicit maps latitude, longitude, and month of the year to a continuous embedding that implicitly captures climatic factors such as temperature and precipitation. To incorporate seasonal trends, we concatenate the embeddings for the months of March, June, September, and December.
@@
Let $\mathbf{X}_{\text{img}}\in\mathbb{R}^{768}$ denote the image embedding obtained from the pretrained DINOv2 encoder, and $\mathbf{X}_{\text{depth}}\in\mathbb{R}^{768}$ denote the depth embedding obtained from the Depth Anything V2 encoder. To incorporate geospatial context, we project the 1024-dimensional embedding produced by Climplicit denoted as $\mathbf{X}_{\text{geo}}\in\mathbb{R}^{1024}$ to a 256-dimensional vector using a trainable linear projection.
The multimodal representation is formed by concatenating all embeddings and is then projected to a 1024-dimensional representation via a linear layer: $\mathbf{Z}=\text{Proj}(\text{concat}(\mathbf{X}_{\text{img}},\mathbf{X}_{\text{depth}},\text{Proj}(\mathbf{X}_{\text{geo}})))\in\mathbb{R}^{1024}.$ The resulting embedding is passed through a residual network of 8 residual blocks with hidden dimension of twice the embedding size. This architecture and embedding dimensions were chosen based on an ablation across multiple configurations (see Appendix). Finally, the output feature representation is passed to four independent heads for trait prediction in our multi-task architecture.
@@
To capture uncertainty in plant trait prediction, each trait-specific prediction head outputs both the predicted value and its associated uncertainty, following the method by (jiang2024uncertainty). For each trait $m\in\{1,\ldots,M\}$, the model predicts two values for each sample $n$: the mean $\hat{\mu}_{n}^{m}$ and the log-scale parameter $\hat{s}_{n}^{m}$, where the scale or standard deviation is given by $b_{n}^{m}=\exp(\hat{s}_{n}^{m})$.
We model the predictive distribution differently for each trait based on its statistical characteristics. For Leaf Area (LA), which exhibits a long-tailed distribution, we use a Laplace distribution parameterized by mean $\hat{\mu}_{n}^{m}$ and scale $b_{n}^{m}=\exp(\hat{s}_{n}^{m})$. The Laplace distribution is more suitable for modeling long-tailed distributions compared to Gaussian distributions  (jiang2024uncertainty).
@@
Citizen science image datasets offer large-scale and diverse data for plant trait modeling but suffer from substantial noise and inconsistencies. Common issues include the presence of non-plant objects, non-representative plant parts, uninformative specimens, and scenes that are too dense, distant, or blurred (Fig. 2). Manual curation at this scale is impractical, and species-level trait annotations often ignore individual variation. Neural networks often memorize noisy labels which harms generalization and makes noise handling essential (lu2022selc; arpit2017closer; zhang2021understanding).  lu2022selc shows models initially learn from clean samples but, past a ‘turning point’, begin memorizing noise, leading to poorer generalization.
@@
Following the approach of previous studies (wolf2022citizen; dechant2024intercomparison), we aggregated the plot-level trait values from sPlotOpen to a 1-degree spatial resolution to generate a global benchmark dataset. We then applied PlantTraitNet to predict trait values using more than 300K globally distributed citizen science observations. The predicted trait values were aggregated to the same 1-degree resolution and filtered to include only grid cells with at least 20 observations, resulting in $\sim$890 grid cells. These predictions were compared against the sPlot-based benchmark dataset using metrices R², mean absolute error (nMAE) normalized by trait range, and Pearson’s correlation (r) on log-transformed values. To account for spatial sampling bias, all metrices are weighted by the area of each 1-degree grid cell. We also compare previously published trait maps against the same sPlotOpen CWM values on overlapping grid cells (boonman2020assessing; butler2017mapping; madani2018future; moreno2018methodology; schiller2021deep; van2014fully; wolf2022citizen). For the ablation study, all models were assessed solely on our validation dataset.
The final model with about $\sim$90M trainable parameters was trained for up to 30 epochs with a batch size of 256 on a single NVIDIA RTX A6000 GPU [using  20GB VRAM]. For selecting the optimal model checkpoint across all traits we compute the Pareto front using the Non-Dominated Sorting (NDS) algorithm (Deb2002NSGAII). We then calculate the hypervolume for all candidate checkpoints on this front (Zitzler2001SPEA2) and select the checkpoint that maximizes the hypervolume. Using (lacoste2019quantifying), we estimate a total of 93.86 kg CO2 emissions for all experiments across seeds. This does not include testing or failed runs and therefore likely underestimates the overall emissions, but it provides a reasonable guideline for future model training.

### v2

@@ Abstract
Extended versionhttps://arxiv.org/abs/2511.06943
@@ Introduction
Terrestrial plants, as the largest primary producers on Earth, contribute about 60% to the global net primary productivity (Field et al. 1998) and play a critical role in the carbon and energy cycles of our Earth system (Pan et al. 2011; Schlesinger and Bernhardt 2020). However, understanding how plants influence these cycles is challenging, as the functioning of plants varies profoundly according to their traits. For instance, traits such as canopy height and leaf area control resource acquisition, while leaf tissue properties, such as nitrogen content or dry matter content, are indicators of plant resilience (Díaz et al. 2016). Although these traits are essential for understanding ecosystem processes, the data on such traits is sparse, as their measurement involves costly field surveys and laboratory analysis. Global plant trait databases such as TRY (Kattge et al. 2011) aggregate thousands of trait measurements from numerous studies and regions, providing an invaluable resource for functional biogeography and ecosystem modeling. However, even with these collective efforts, significant gaps persist in the geographic coverage of trait data across biomes, ecosystems, and species, constraining our ability to fully understand and predict global patterns of vegetation function and change  (Díaz et al. 2016; Kattge et al. 2020).
@@
Given the strong link between plant morphology and function, plant photographs in concert with computer vision offer a promising avenue for large-scale estimation of plant traits. Citizen science platforms such as iNaturalist (Su and Maji 2021) and Pl@ntNet (Garcin et al. 2021) have collected more than 50 million research-grade plant photographs around the world, creating a unique resource for uncovering global plant trait distributions (Wolf et al. 2022). These datasets, primarily curated for the identification of plant species, provide plant images and species labels, but do not provide trait annotations  (Goëau et al. 2025; Stevens et al. 2024; Van Horn et al. 2018). However, prior work has demonstrated that trait information for these species can be indirectly obtained by linking species names from citizen science records to trait databases such as TRY (Schiller et al. 2021; Wolf et al. 2022). Through this species-level matching, trait values can be weakly assigned to images, enabling the construction of large-scale, trait-annotated image datasets. These datasets can then be used to train scalable computer vision models for trait prediction from images (Schiller et al. 2021). Such models enable direct trait estimation from photographs, independent of whether the species is known or if a record exists in a trait database. Here, we attempt to advance this approach by predicting multiple traits simultaneously, leveraging shared visual features and underlying trait correlations. Subsequently, we spatially aggregate trait predictions derived from individual geotagged photographs to create global, gridded geospatial maps representing the trait distributions across plant communities and ecosystems  (Schiller et al. 2021).
@@
The geolocation of each photograph not only enables the spatial aggregation of the predictions into geospatial maps but also allows for the integration of spatial context into the prediction process itself  (Schiller et al. 2021). For instance, climate, including temperature and precipitation, or phenological information from satellite data is known to be the key variable shaping global trait distributions, making it a promising predictor  (Bruelheide et al. 2018; Schiller et al. 2021; Joswig et al. 2022). However, integrating large-scale geospatial products can be challenging due to data gaps and feature selection (Lusk et al. 2025).
@@
Recent advances in geospatial foundation models (GeoFMs) now support the seamless integration of such context into downstream tasks. Examples include Climplicit (Dollinger et al. 2025), which encodes climate information, and SatCLIP (Klemmer et al. 2025), which leverages satellite Earth observation data. Such GeoFMs have demonstrated strong generalization in global mapping applications. In this study, we test the integration of such GeoFMs into visual trait prediction to enhance performance through geospatial context.
@@
An important challenge for computer vision with citizen science data is its inherent heterogeneity and noise (Sierra et al. 2024), ranging from inconsistent image quality due to varied photo acquisition methods (feature noise) to ambiguous trait annotations from weak supervision (label noise). Such data characteristics may result in implausible predictions and leave an imprint on the aggregation into global trait products. Moreover, such training data noise can substantially degrade a model’s ability to generalize to unseen data  (Lu and He 2022; Arpit et al. 2017; Zhang et al. 2021). To overcome both feature and label noise in the citizen-science data, we propose an uncertainty-aware probabilistic deep learning framework that estimates predictive uncertainty. The predicted uncertainty is used to dynamically down-weight highly noisy samples and to filter out unreliable data points, thereby reducing overfitting to spurious patterns.
@@
  • We apply PlantTraitNet on more than 300K independent samples of citizen science photos and spatially aggregate the predictions to global trait maps. A benchmark against globally distributed vegetation survey data (sPlotOpen) revealed that these PlantTraitNet-derived traits maps consistently outperform previous global trait products.
@@
Pioneering work by Schiller et al. (2021) showed that plant traits, such as height, nitrogen content, specific leaf area, or leaf area, can be predicted from citizen science images using weak supervision, where species-level trait labels are derived from the TRY database (Kattge et al. 2020). While  Schiller et al. (2021) focused on single-task models, Cherif et al. (2023) showed that predicting multiple plant traits simultaneously can exploit trait-trait correlations and shared features in the predictor data.
However, Schiller et al. (2021) did not assess whether weak supervision enables capturing within-species trait variation (e.g., size differences among individuals of the same species). Moreover,  Schiller et al. (2021) did not test how aggregating individual predictions on a global scale resembles large-scale trait variation across the biosphere. Wolf et al. (2022) provided an approach to validate global trait maps using vegetation survey data of plant communities from the collaborative initiative sPlot (Bruelheide et al. 2018; Sabatini et al. 2021) linked with trait data from the TRY database (Kattge et al. 2020). This approach provides an effective means to evaluate the potential of computer vision models for generating trait maps at global scale.
A persistent challenge with citizen science data is the noise in both images and labels (Sierra et al. 2024; Schiller et al. 2021), often structured spatially. Such noise can bias both inference and training, as deep networks tend to memorize noisy labels, compromising generalization (Lu and He 2022; Arpit et al. 2017; Zhang et al. 2021).
@@
  • Using visual and depth-based foundation models (Oquab et al. 2023; Yang et al. 2024) to better represent heterogeneous plant imagery.
  • Leveraging multi-task learning to exploit trait correlations (Cherif et al. 2023).
  • Incorporating uncertainty-aware training to address label noise (Yeo et al. 2021; Jiang et al. 2024).
  • Benchmarking global trait predictions against sPlot vegetation survey data (Wolf et al. 2022).
@@
Figure 1: Geographic coverage of the citizen science data (top) and independent benchmark reference data (bottom) from vegetation surveys (sPlotOpen, Sabatini et al. 2021).
@@
To predict plant traits at a global scale, we utilize two large-scale citizen science datasets: iNaturalist (GBIF.org 2025; Su and Maji 2021) and Pl@ntNet-300K (Garcin et al. 2021). These datasets consist of plant images annotated with species labels and geolocations but lack direct trait measurements. Following Schiller et al. (2021), we weakly annotate each image using species-level trait distributions from the TRY database (Kattge et al. 2020), based on the premise that interspecific trait variation (variation between species) generally exceeds intraspecific variation (variation within species) (Dong et al. 2020; Wright et al. 2017).
We model each trait as a normal distribution per species, using TRY-derived means and standard deviations, and sample trait values within the interquartile range to reduce outlier influence. To account for intraspecific variability, we resample traits for each image at every training epoch  (Schiller et al. 2021).
@@
We further reduce noise through model-driven uncertainty estimates (see Methodology). The final dataset includes approximately 220K training images across 5K species and over 80K images in the validation set.
@@
For evaluation, we use the sPlotOpen database (Sabatini et al. 2021). The georeferenced sPlotOpen records represent plant community compositions, which were linked with trait data from the TRY database (Kattge et al. 2020). This data provides global trait maps of community-weighted mean (CWM) trait values.
@@
Figure 4: Overview of the pipeline. We filter weakly labeled citizen science data (Raw data) based on high model uncertainty (Step 1) and large residuals from species trait medians (Step 2). We use this refined data for training the models (Step 3), which are evaluated by comparing spatially aggregated predictions (1° resolution) against overlapping vegetation surveys (sPlotOpen).
@@
Given an input image $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$, we use the pretrained DINOv2 ViT-B/14 encoder to extract a sequence of patch-level feature embeddings $\in\mathbb{R}^{N\times C}$  (Oquab et al. 2023).
@@
A novel addition to our architecture is the use of depth priors from foundation models for monocular depth estimation. While standard 2D RGB images lack explicit three-dimensional spatial cues, depth information encodes the distance between the sensor and surface points on the plant, enabling a more accurate reconstruction of plant morphology and structure. To incorporate depth, we use the pretrained and frozen encoder from the Depth-Anything-V2 (DA-V2) model  (Yang et al. 2024), denoted as $h$. Although various models could be used, we adopt DA-V2 for its strong generalization capabilities, attributed to its training on large-scale labeled and unlabeled datasets and its student-teacher distillation framework. We use the ViT-B variant, which outputs a set of embeddings $h(I)\in\mathbb{R}^{N\times C}$. Similar to the image encoder, we apply adaptive average pooling along the patch token dimension $N$, reducing it to a 64-dimensional representation. This pooled output is flattened and passed through a MLP to generate the depth prior embedding of dimension 768.
@@
Plants are tailored to local climatic conditions, such as precipitation and temperature, through their traits (Joswig et al. 2022).
To incorporate this climatic context as a cue in the prediction process, we integrate Climplicit (Dollinger et al. 2025) into our architecture, a spatio-temporal geo-location encoder trained on the CHELSA climate dataset (Karger et al. 2017). Climplicit maps latitude, longitude, and month of the year to a continuous embedding that implicitly captures climatic factors such as temperature and precipitation. To incorporate seasonal trends, we concatenate the embeddings for the months of March, June, September, and December.
@@
Let $\mathbf{X}_{\text{img}}\in\mathbb{R}^{768}$ denote the image embedding obtained from the pretrained DINOv2 encoder, and $\mathbf{X}_{\text{depth}}\in\mathbb{R}^{768}$ denote the depth embedding obtained from the DA-V2 encoder. To incorporate geospatial context, we project the 1024-dimensional embedding produced by Climplicit denoted as $\mathbf{X}_{\text{geo}}\in\mathbb{R}^{1024}$ to a 256-dimensional vector using a trainable linear projection.
The multimodal representation is formed by concatenating all embeddings and is then projected to a 1024-dimensional representation via a linear layer: $\mathbf{Z}=\text{Proj}(\text{concat}(\mathbf{X}_{\text{img}},\mathbf{X}_{\text{depth}},\text{Proj}(\mathbf{X}_{\text{geo}})))\in\mathbb{R}^{1024}.$ The resulting embedding is passed through a residual network of 8 residual blocks with hidden dimension twice the embedding size. This architecture and embedding dimensions were chosen based on an ablation across multiple configurations (see Appendix). Finally, the output feature representation is passed to four independent heads for trait prediction in our multi-task architecture.
@@
To capture uncertainty in plant trait prediction, each trait-specific prediction head outputs both the predicted value and its associated uncertainty, following the method by (Jiang et al. 2024). For each trait $m\in\{1,\ldots,M\}$, the model predicts two values for each sample $n$: the mean $\hat{\mu}_{n}^{m}$ and the log-scale parameter $\hat{s}_{n}^{m}$, where the scale or standard deviation is given by $b_{n}^{m}=\exp(\hat{s}_{n}^{m})$.
We model the predictive distribution differently for each trait based on its statistical characteristics. For Leaf Area (LA), which exhibits a long-tailed distribution, we use a Laplace distribution parameterized by mean $\hat{\mu}_{n}^{m}$ and scale $b_{n}^{m}=\exp(\hat{s}_{n}^{m})$. The Laplace distribution is more suitable for modeling long-tailed distributions compared to Gaussian distributions  (Jiang et al. 2024).
@@
Citizen science image datasets offer large-scale and diverse data for plant trait modeling but suffer from substantial noise and inconsistencies. Common issues include the presence of non-plant objects, non-representative plant parts, uninformative specimens, and scenes that are too dense, distant, or blurred (Fig. 2). Manual curation at this scale is impractical, and species-level trait annotations often ignore individual variation. Neural networks often memorize noisy labels which harms generalization and makes noise handling essential (Lu and He 2022; Arpit et al. 2017; Zhang et al. 2021).  Lu and He (2022) shows models initially learn from clean samples but, past a ‘turning point’, begin memorizing noise, leading to poorer generalization.
@@
Following the approach of previous studies (Wolf et al. 2022; Dechant et al. 2024), we aggregated the plot-level trait values from sPlotOpen to a 1-degree spatial resolution to generate a global benchmark dataset. We then applied PlantTraitNet to predict trait values using more than 300K globally distributed citizen science observations. The predicted trait values were aggregated to the same 1-degree resolution and filtered to include only grid cells with at least 20 observations, resulting in over 890 grid cells. To ensure a robust and interpretable evaluation, we used complementary metrics that capture different aspects of model performance: the coefficient of determination (R²) quantifies the proportion of variance in observed trait values explained by the predictions, the normalized mean absolute error (nMAE) measures the average deviation normalized by the observed trait range, and Pearson’s correlation coefficient (r) on log-transformed values quantifies the linear relationship between predicted and observed data. To account for spatial sampling bias, all metrics are weighted by the area of each 1-degree grid cell. We also compare previously published trait maps against the same sPlotOpen CWM values on overlapping grid cells (Boonman et al. 2020; Butler et al. 2017; Madani et al. 2018; Moreno-Martínez et al. 2018; Schiller et al. 2021; Van Bodegom et al. 2014; Wolf et al. 2022). For the ablation study, all models were assessed solely on our validation dataset.
The final model, with approximately 90M trainable parameters, was trained for up to 30 epochs with a batch size of 256 on a single NVIDIA RTX A6000 GPU (using approximately 20 GB VRAM). To select the optimal model checkpoint across all traits, we compute the Pareto front using the Non-Dominated Sorting (NDS) algorithm (Deb et al. 2002). We then calculate the hypervolume for all candidate checkpoints on this front (Zitzler et al. 2001) and select the checkpoint that maximizes the hypervolume. Following Lacoste et al. (2019), we estimate a total of 93.86 kg CO2 emissions for all experiments across seeds. This estimate excludes testing and failed runs and therefore likely underestimates the total emissions, but it provides a reasonable guideline for future model training.